# WOMD-Reasoning: A Large-Scale Language Dataset for Interaction and Driving Intentions Reasoning

### v1

@@ Abstract
We propose Waymo Open Motion Dataset-Reasoning (WOMD-Reasoning), a language annotation dataset built on WOMD***Waymo Open Motion Dataset and WOMD are trademarks of Waymo LLC, and are used here by permission., with a focus on describing and reasoning interactions and intentions in driving scenarios. Previous language datasets primarily captured interactions caused by close distances. However, interactions induced by traffic rules and human intentions, which can occur over long distances, are yet sufficiently covered, despite being very common and more challenging for prediction or planning models to understand. Therefore, our WOMD-Reasoning focuses extensively on these interactions, providing a total of 409k Q&As for varying types of interactions. Additionally, WOMD-Reasoning presents by far the largest Q&A dataset on real-world driving scenarios, with around 3 million Q&As covering various topics of autonomous driving from map descriptions, motion status descriptions, to narratives and analyses of agents’ interactions, behaviors, and intentions. This extensive textual information enables fine-tuning driving-related Large Language Models (LLMs) for a wide range of applications like scene description, prediction, planning, etc. By incorporating interaction and intention language from WOMD-Reasoning, we see significant enhancements in the performance of the state-of-the-art trajectory prediction model, Multipath++, with improvements of 10.14% in $MR_{6}$ and 6.90% in $minFDE_{6}$, proving the effectiveness of WOMD-Reasoning. We hope WOMD-Reasoning would empower LLMs in driving to offer better interaction understanding and behavioral reasoning. The dataset is available on https://waymo.com/open/download.
Figure 1: A typical example of interaction descriptions and reasoning in WOMD-Reasoning. In this case, WOMD-Reasoning successfully captures the traffic rule-induced interaction between the ego agent and agent # 0, attributing it correctly to the stop signs.
@@ 1 Introduction
Large-guided driving datasets and models have drawn much attention in autonomous driving. For example, methods based on large language models (LLMs) [27, 1] have shown effectiveness in improving explainability, controllability, and performance of driving tasks [25, 5, 10, 14, 33]. Various language datasets [25, 19, 12, 23, 14, 11] have been proposed for fine-tuning these LLMs. However, due to the complexity and variability of driving scenarios and driver behaviors, existing datasets often focus on specific scenes without covering critical cases, e.g., reasoning about interactions and human intentions, which is vital for the advancement of complex decision-making processes in autonomous driving systems [29, 36, 37]. For example, in BDD-X [12], interactions are often limited in spatial relationships, e.g., ‘The car moves back into the left lane because the school bus in front of it is stopping.’, while in an example, DriveLM [25] attributes interaction to ‘keep a safe distance’ without further detailed explanations. As a result, most existing works have drawbacks, either insufficient in quantity or only covering interactions induced by proximity, as shown in our analysis in Table 1.
@@
Considering the critical importance and difficulty of interaction modeling for autonomous driving [8, 21], it is essential to incorporate comprehensive interactive and behavioral reasoning for language-based driving. Real-world interactions are often driven by human knowledge and behavior, such as traffic rules, signals, and intentions like overtaking. However, these interactions are rarely covered in previous datasets, leading to suboptimal performance of LLMs fine-tuned on these datasets. Additionally, most existing language datasets are not large enough to support LLM fine-tuning across diverse tasks such as scene description, prediction, and planning.
@@
Therefore, we build WOMD-Reasoning, a large-scale language Q&A dataset based on WOMD [9], with a focus on interaction descriptions and reasoning. To incorporate interactions induced by traffic rules and human intentions, we build an automated data curation pipeline based on prompting ChatGPT [1, 18] with a set of well-designed prompts. Together with a rule-based translator to convert motion data into language, we create an automated pipeline for generating the language dataset, achieving an overall accuracy rate of over 90%. We then manually verify and correct a subset to provide a high-quality set suitable for supervised fine-tuning (SFT) with LLMs. To the best of our knowledge, WOMD-Reasoning is the currently largest real-world language dataset for driving.
The resultant WOMD-Reasoning dataset successfully covers interactions induced by traffic rules and human intentions with 409k Q&As, like those shown in Figure 1. As indicated in Table 1, it contains the largest number of total Q&As and interaction-specific Q&As, supporting a wide range of applications including scene description, prediction, and planning. We empirically find that language in WOMD-Reasoning can significantly enhance the performance of vehicle trajectory prediction tasks. Qualitative results indicate a good consistency between these performance improvements and the corresponding language inputs. Our contributions can be summarized in the following:
@@
  • We propose WOMD-Reasoning, a language dataset centered on interaction descriptions and reasoning. It provides extensive insights into critical but previously overlooked interactions induced by traffic rules and human intentions.
  • We develop an automatic language labeling pipeline, leveraging a rule-based translator to interpret motion data into language descriptions, and a set of manual prompts for ChatGPT to generate Q&A pairs. As a result, WOMD-Reasoning emerges as, to the best of our knowledge, the largest language dataset for real-world driving, comprising approximately 3 million Q&A pairs supporting scene description, prediction, and planning tasks.
  • We validate the effectiveness of WOMD-Reasoning in vehicle trajectory prediction tasks. By training Multipath++ on WOMD+WOMD-Reasoning, significant performance improvements have been observed, e.g., 10.14% in $MR_{6}$ and 6.90% in $minADE_{6}$. Our findings suggest a notable correlation between the language input and the improvements in prediction accuracy, proving the effectiveness of interaction narratives in WOMD-Reasoning.
@@
Table 1: Comparison between WOMD-Reasoning and Previous Real-world Language Datasets for Driving Scenarios. Compared to previous datasets, WOMD-Reasoning covers comprehensive kinds of interactions, with a significantly higher size supporting a wide range of potential applications
@@ 2 Related Works
Application of Language Modal in Autonomous Driving. Language is widely utilized in autonomous driving models, as incorporating language into implicit features could enhance the explain-ability [34, 22, 26, 3, 15]. For example, Kuo et al. [13] employ language features in the trajectory prediction model, ensuring both explainability and consistency between language and predictions. Similarly, Zhong et al. [38] use language input and large language models (LLMs) to generate conditions in diffusion models, thereby controlling the generation of driving scenarios. Recently, general driving agents have been developed [33, 5, 25, 16], aiming to integrate all functions—perception, prediction, and planning—into a single language agent. However, we observe that their performance in interaction analysis is not entirely satisfying. This shortcoming is largely attributed to the lack of interaction analysis in language datasets specific to driving scenarios.
@@
Language Datasets in Autonomous Driving. Language datasets for driving scenarios have recently been developed to support related LLM-based work. Due to the controllability of simulations, several simulation-based datasets have been created [5, 25]. However, simulations often fail to capture real-world interactions. To address this, recent studies [12, 23, 14] incorporate human labeling of real-world driving scenarios to ensure the inclusion of genuine interactions. Nevertheless, obtaining high-quality human labels is labor-intensive and costly, limiting the size and coverage of these datasets. Moreover, interactions driven by traffic rules and human intentions have often been overlooked. To streamline the labeling process, some studies [7, 17, 19, 25] have generated rule-based labels, although these typically cover only basic language elements like scene descriptions, leaving interaction analysis to human labelers. To minimize human labor, we first employ manual rules and then utilize ChatGPT-4 [18] to build our dataset automatically. The resulting dataset, WOMD-Reasoning, is rich in interaction details, particularly those stemming from traffic rules and human intentions. The designed automated data-curation pipeline enables the dataset to be significantly larger. We also create a high-quality human-verified subset that can be used for supervised fine-tuning. Preliminary human evaluation provides positive feedback with an overall accuracy of over 90%.
@@ 3 Method
Building language datasets would involve intense human labor, which is one of the main reasons why previous datasets have insufficient interaction analysis and a very limited overall size [24]. To reduce human labor while obtaining reasonably useful data on vehicle interactions, we propose a fully automatic data-curation pipeline to label the motion dataset with language: We first develop a rule-based program to interpret the motion dataset, which contains trajectories and the HD map, into language; and then build a set of prompts to utilize the reasoning abilities of ChatGPT-4 [18] to generate interaction analysis and reasoning and organize the results into the target Q&A format.
3.1 Automatic Translation of Driving Scenarios into Language Descriptions
To ensure enough information for analyzing interactions involving the ego agent, we translate the motion data into an ego-centric description of the traffic scenario. Specifically, before describing the ego agent status, we first narrate the map environment around the ego agent, including information about related intersections, lanes, stop signs, crosswalks, speed bumps, etc. The above information provides a comprehensive spatial guidance that facilitates describing the motion of either ego or other agents accurately. Then, we describe the status of the ego agent in the scenario, including its motion status (e.g., the velocity, acceleration, etc.), and the positional status (e.g., the lane it occupies and its position related to the intersection center). Furthermore, its related traffic light and signal information is also summarized in the description. Besides the ego agent, the depiction of other agents, including other vehicles, bicycles, and pedestrians is also crucial for GPT to analyze the interactions. Therefore, we then describe every other agent in the scenario in the same way as we describe the ego agent. We additionally include the positional relation between the other agent and the ego one to support the identification of every agent’s position.
Utilizing the methods above, we are able to translate motion datasets into languages. Due to the limitations of computational resources of GPT, we only choose to translate those scenarios that have ”objects_of_interest” labels in WOMD, which indicates confirmed significant interactions happening in the scenario. We build two subsets of WOMD-Reasoning with the same setting: the training set is built on the training set of WOMD, while the validation set is built on the interactive validation set of WOMD. In total, we translate 63k scenarios into language. An example of the translation can be found in Table 5 in the Appendix.
3.2 GPT-based Interaction Analysis
With the language translation of WOMD as input, we further design a set of prompts to utilize ChatGPT-4 [18] to build the Q&A dataset, taking advantage of its abilities in analyzing interactions. The prompts come with 4 main parts. (1) System Prompt, which lets the GPT know its role and the format of its input, i.e. the language translation. (2) Responsibility Prompt, which describes the responsibilities of the GPT, including the questions it should ask and answer, as well as the output format it should use. This will be talked about in detail in the next paragraph. (3) Rules Prompt, which contains global rules to guide the output and the analysis. (4) In-context Prompt, which provides a few human-written examples of input-output pairs to perform in-context learning. An example of each part of the prompts can be found in Table 6, 7, 8, and 9 respectively in the Appendix.
The questions in the responsibility part of the prompt decide what would be present in the generated language Q&A dataset. To provide a comprehensive language dataset, and to present interaction analysis as accurately as possible, we organize the questions into a chain to enable a well-designed chain-of-thought [30, 31, 6]: First, we ask questions about the map environment, the ego and other agents’ motion. These questions would let the GPT provide Q&As useful for fine-tuning scene description agents while enhancing the GPT’s understanding of the input information at the same time. Then, we ask questions about the interactions that occur in the observable future period. To include traffic rule-induced and intention-induced interactions, we ask the GPT to think about the interaction between each pair of agents by answering a sequence of questions: Q1: Are the two agents vehicles, and are they close to each other with traffic rules or signals governing their movements? If the answer is yes, we request GPT to give an analysis of their yielding relations according to the traffic rules. Otherwise, we ask Q2: Are the two agents vehicle and pedestrian respectively with intention conflicts? If so, the vehicle generally should yield to the pedestrian for safety. If both questions do not help find the interaction relations, we come to Q3: Does the scene show patterns of certain intention-induced interaction? We provide descriptions of a few common interactions like overtaking, following, etc. GPT can choose the best fit from these patterns to try to cover more interaction types. Based on this pipeline, we are able to provide interaction analysis with rich information on traffic rule-induced and human intention-induced interactions. And after that, finally, we are able to wrap up the interactions between the ego and all other agents, to provide the intention of the ego agent.
@@ 3.2 GPT-based Interaction Analysis
Our analysis process is performed on Microsoft Azure ChatGPT-4-Turbo API, which costs around 12,750 USD. The details of WOMD-Reasoning will be provided in the following paragraph.
Figure 2: An Additional Example of Traffic Rule-based Interaction and Intention in WOMD-Reasoning. In this scenario, WOMD-Reasoning completely covers the traffic light-controlled yielding interaction between the ego agent and agent # 1.

### v2

# WOMD-Reasoning: A Large-Scale Dataset and Benchmark for Interaction and Intention Reasoning in Driving

@@ Abstract
We propose Waymo Open Motion Dataset-Reasoning (WOMD-Reasoning), a comprehensive large-scale dataset with 3 million Q&As built on WOMD focusing on describing and reasoning interactions and intentions in driving scenarios. Existing language datasets for driving primarily capture interactions caused by close distances. However, interactions induced by traffic rules and human intentions, which can occur over long distances, are yet sufficiently covered. To address this, WOMD-Reasoning presents by far the largest multi-modal Q&A dataset on real-world driving scenarios, covering a wide range of driving topics from map descriptions and motion status descriptions to narratives and analyses of agents’ interactions, behaviors, and intentions. We further introduce Motion-LLaVA, a motion-language model fine-tuned on the proposed dataset with robust interaction reasoning capabilities. We benchmark its performance across various configurations including different input modalities, reasoning techniques, and network architectures. The robust, diverse, and multi-modal nature of WOMD-Reasoning highlights its potential to advance future autonomous driving research and enable a broad range of applications. The dataset and its vision modal extension are available at https://waymo.com/open/download, and the codes & prompts to build it are available at https://github.com/yhli123/WOMD-Reasoning.
Figure 1: Two examples of interaction descriptions and reasoning in WOMD-Reasoning dataset. (a) This example captures the traffic rule-induced interaction between the ego agent and agent #0, attributing it correctly to the stop signs. (b) This example shows the traffic light-controlled yielding interaction between the ego agent and agent #1. The front-view visualization is created by importing the traffic scenario into MetaDrive simulator [24].
@@ 1 Introduction
Language-guided driving datasets and models have drawn much attention in autonomous driving. For example, methods based on large language models (LLMs) [50, 4] improve explainability, controllability, and performance of driving tasks [7, 14, 30, 57, 23, 2, 36, 26, 9, 29, 60, 16, 11, 17, 53].
@@
Various language-based datasets [45, 37, 20, 41, 30, 18, 65, 48, 2, 36, 26, 29] have been proposed to fine-tune these LLMs. However, due to the complexity and variability of driving scenarios and driver behaviors, existing datasets often focus on describing specific scenes without reasoning about the interactions, like traffic rules and human intentions behind them, which are vital for the safety-critical decision-making in autonomous driving systems [54, 62, 63]. For example, in BDD-X [20], interactions are often limited in spatial relationships, e.g., ‘The car moves back into the left lane because the school bus in front of it is stopping.’; while in another example, DriveLM [45] attributes interaction to ‘keep a safe distance’ without further detailed explanations. As a result, most existing works are either insufficient in quantity or only cover interactions induced by proximity, as shown in our analysis in Table 1.
@@
Considering the critical importance and difficulty of interaction modeling for autonomous driving [12, 39], it is essential to incorporate comprehensive interactive and behavioral reasoning for language-based driving. Real-world interactions are often driven by human knowledge and behavior, such as traffic rules, signals, and intentions like overtaking. However, these interactions are rarely included in previous datasets, leading to suboptimal performance of LLMs fine-tuned on these datasets. Additionally, most datasets in driving are not large enough to support multi-modal fine-tuning across diverse tasks such as scene description, prediction, and planning.
@@
This work introduces WOMD-Reasoning dataset, a large-scale multi-modal dataset centered on language Q&As based on WOMD [13], with a specific focus on interaction descriptions and reasoning. To incorporate interactions induced by traffic rules and human intentions, we build an automated data curation pipeline by prompting ChatGPT [4, 34] with a set of well-designed prompts. Together with a rule-based translator to convert motion data into language, we create an automated pipeline for generating the language dataset and achieve an approximate accuracy rate of over 90%. We then manually verify and correct a subset to provide a high-quality set suitable for supervised fine-tuning (SFT) with LLMs. To the best of our knowledge, WOMD-Reasoning is currently the largest real-world language dataset for driving. WOMD-Reasoning dataset covers interactions induced by traffic rules and human intentions with 409k Q&As, like those shown in Figure 1. As indicated in Table 1, it contains the largest number of total Q&As and interaction-specific Q&As, supporting a wide range of applications including scene description, prediction, and planning. Additionally, to facilitate multi-modal learning across motion, language, and vision modalities, for each motion data, we render both the corresponding bird’s eye view and front view images by the MetaDrive simulator [24].
Based on WOMD-Reasoning, we build our Motion-LLaVA upon LLaVA [28], which can accurately answer driving-related questions. To comprehensively absorb detailed motion information, we further replace LLaVA’s vision encoder with a motion vector encoder [51], a motion prediction model. With this design and a set of effective training and inference strategies ablated, Motion-LLaVA achieves high accuracy in both driving scenario understanding and interaction analysis, outperforming baselines by large margins. We further benchmark several settings including multi-modal model fine-tuning strategies on WOMD-Reasoning to show the advantage of Motion-LLaVA. It unleashes the analytic abilities of motion language models to provide in-depth descriptions and reasoning on driving scenarios.
Our contributions can be summarized as follows:
@@
  • We propose WOMD-Reasoning, the largest multi-modal dataset with 3 million Q&A pairs centered on interaction and intention reasoning in driving. It provides extensive insights into critical but previously overlooked interactions induced by traffic rules and human intentions.
  • We introduce Motion-LLaVA, a motion-language model designed for WOMD-Reasoning with robust capabilities, by benchmarking task performance across various configurations, including varying input modalities (vision v.s. motion), reasoning techniques (single step v.s. chain-of-thought), and network architectures (motion encoder pre-training).
  • Fine-tuned with WOMD-Reasoning, Motion-LLaVA is capable of providing detailed and insightful descriptions and interaction reasonings on various driving scenarios. Extensive evaluations with various metrics validate its wide coverage and high quality, especially in reasoning traffic rule-induced interactions.
@@
Table 1: Comparison between WOMD-Reasoning Dataset and Previous Real-world Language Datasets for Driving Scenarios. Compared to previous datasets, WOMD-Reasoning covers comprehensive kinds of interactions, with a significantly higher size supporting a wide range of potential applications.
@@ 2 Related Work
Application of Language Modal in Autonomous Driving. In recent years, language has been widely explored in autonomous driving, as incorporating language into implicit features could enhance the explain-ability [59, 40, 46, 6, 31]. For example, Kuo et al. [22] employ language features in the trajectory prediction model, ensuring both explainability and consistency between language and predictions. Similarly, Zhong et al. [64] use language input and LLMs to generate conditions in diffusion models, thereby controlling the generation of driving scenarios. Recently, general driving agents have been developed [57, 7, 45, 32], aiming to integrate all functions—perception, prediction, and planning—into a single language agent. However, we observe that their performance in interaction analysis is not entirely satisfying. This shortcoming is largely attributed to the lack of interaction analysis in language datasets specific to driving scenarios.
@@
Language Datasets in Autonomous Driving. Language datasets for driving scenarios have recently been developed to support related LLM-based work. Due to the controllability of simulations, several simulation-based datasets have been created [7, 45]. However, simulations often fail to capture real-world interactions. To address this, recent studies [20, 41, 30] incorporate human labeling of real-world driving scenarios to ensure the inclusion of genuine interactions. Nevertheless, obtaining high-quality human labels is labor-intensive and costly, limiting the size and coverage of these datasets. Moreover, interactions driven by traffic rules and human intentions have often been overlooked. To streamline the labeling process, some studies [10, 33, 37, 45] have generated rule-based labels, although these typically cover only basic language elements like scene descriptions, leaving interaction analysis to human labelers. To minimize human labor, we first employ manual rules and then utilize ChatGPT-4 [34] to build our dataset automatically. The resulting dataset, WOMD-Reasoning, is rich in interaction details, particularly those stemming from traffic rules and human intentions. The designed automated data-curation pipeline enables the dataset to be significantly larger. We also create a high-quality human-verified subset. Preliminary human evaluation provides positive feedback with an approximate accuracy of 90%.
Fine-tuning LLM for Driving Tasks. Recent years have witnessed a surge of research on leveraging LLMs and vision-language models (VLMs) for driving applications [43, 7, 19, 58, 45, 44, 31, 47]. However, handling heterogeneous inputs such as vectorized motion representation remains challenging due to their domain gaps to most existing VLMs and LLMs. Notable prior work, such as [31], designed natural language prompt structures to accommodate numerical motion data, while another approach [7] proposed training LLMs with custom-built motion vector encoders.
Our Motion-LLaVA approach is inspired by the design of [7], augmenting existing LLM architectures with tailored components. In our study, we scale the training dataset by approximately 20 times and use real-world scenarios instead of simulation-based data facilitated by WOMD-Reasoning. Furthermore, we employ a motion prediction model as a motion vector encoder, utilizing its prior knowledge to streamline the training pipeline. To enhance reasoning accuracy, we adopt a Chain-of-Thought (CoT) [55] approach, allowing the model to base its reasoning on factual grounding, mitigating potential hallucinations. With WOMD-Reasoning, our Motion-LLaVA demonstrates the practicality of applying VLMs to real-world vectorized motion data to accurately answer diverse driving-related questions correctly.
@@ 3.1 Building WOMD-Reasoning Dataset
Manually constructing autonomous driving datasets with fine-grained natural-language labels regarding the interactions among road participants would involve intense human labor, which is one of the main reasons why previous datasets have insufficient interaction analysis and a very limited overall size [42]. To reduce human labor while obtaining reasonably useful data on vehicle interactions, we propose a fully automatic data-curation pipeline to label the WOMD dataset with language: We first develop a rule-based program to interpret the motion dataset, which contains trajectories and the HD map, into language; and then build a set of prompts to utilize the reasoning abilities of ChatGPT-4 [34] to generate interaction analysis and reasoning, and organize the results into the target Q&A format. Details of each step are attached in the Appendix A.1. Codes for interpreting the motion dataset as well as the full prompts for ChatGPT-4 are available at https://github.com/yhli123/WOMD-Reasoning.
@@
Our analysis process is performed on Microsoft Azure ChatGPT-4-Turbo API, which costs around 12,750 USD. The details of WOMD-Reasoning will be provided in Part 4.
3.2 Motion-LLaVA: an Multi-modal Fine-tuning Method on WOMD-Reasoning
To showcase the application of WOMD-Reasoning dataset, we perform a set of multi-modal fine-tuning experiments on it to build our Motion-LLaVA. LLaVA [28] is chosen as the pre-trained multi-modal model, in order to accommodate motion data input and text prompts input at the same time, where the vision encoder should be adapted to accommodate motion data inputs. Previous work [7] directly attended vectorized motion data with learnable queries to aggregate motion information. Inspired by [44], which trained its vision encoder with future trajectories as supervision, we utilize encoders from motion prediction models for encoding motion data. Specifically in our work, we adopted the encoder component of the motion prediction model MultiPath++ [51] as the motion vector encoder for the LLM. Ablations in Table 5 prove the helpfulness of knowledge prior in this encoder for fine-tuning. We also ablate some detailed fine-tuning strategies in Appendix A.7. Our Motion-LLaVA fine-tuning pipeline, summarized as the Motion-LLaVA method, is illustrated in Figure 3.
During training, we unfreeze all components including the motion vector encoder, and train on all fields of Q&A pairs in WOMD-Reasoning simultaneously to avoid the potential catastrophic forgetting. To prevent Motion-LLaVA from forgetting its inherent LLM’s knowledge, we mix up a few questions irrelevant to the current driving scenario into the training process. [61] During inference, we adopt Chain-of-Thought (CoT) Reasoning strategy [55] to reduce the hallucination. Specifically, we let Motion-LLaVA first answer the facts questions, which include questions in map environment, ego, and other agents’ motion status. Then, we aggregate these answers and feed them into Motion-LLaVA as the contexts for answering reasoning questions, which refers to questions in interactions and intentions analysis. Examples of these contexts can be found in the Appendix A.5. We ablate in Part 5.4 that such a CoT strategy can effectively improve the quality of answering reasoning questions.
4 WOMD-Reasoning Dataset
In this section, we provide detailed information about WOMD-Reasoning. We first present the organization and statistics of the dataset, then provide quality analysis based on case studies and preliminary human evaluations. Finally, we compare our dataset to previous ones to show the strengths of WOMD-Reasoning in diverse interactions, huge size, and wide potential applications.
4.1 Dataset Overview
Table 2 provides the statistics of Q&As in WOMD-Reasoning dataset, as well as in every Q&A class. Our Q&As can be classified into 5 categories: The first 3 categories are Factual contents: Map Environment, which includes the existence and category of intersections, existence and places of stop signs and crosswalks, and count of lanes, etc. Ego Agent’s Motion Status and Other (Surrounding) Agents’ Motion status, which describe each agent’s speed, acceleration, direction, related traffic light and sign, and relative position to the intersection center as well as to the ego agent. The other 2 categories are Reasoning contents: Interaction, which includes a summary of interactions between each surrounding agent and the ego agent and the reasoning behind such interactions. Intention, which sums up all interaction information to give the predicted intention of the ego agent, considering its responses to all the interactions. The factual description of the map environment and agents’ motion status in the dataset are good for training language scene describers, while the interaction and intention reasoning parts can be used for training language-based prediction and planning agents.
We claim that to the best of our knowledge, WOMD-Reasoning is the largest real-world language dataset for driving. As shown in Table 1, we provide roughly 3 million Q&A pairs for 63k scenes in WOMD, which are 6-113 times greater than those of other datasets we compare to. Furthermore, WOMD-Reasoning contains 346k unique traffic rule-induced and intention-induced interaction Q&As, as well as 63k comprehensive intention predictions of the ego agent containing responses to the interactions, which rarely exist in previous datasets. We will show these features through examples and statistics in the following section.